Race One Motors is an Indonesian car manufacturer. At its largest manufacturing facility, in Jakarta, the company produces subcomponents at a rate of 300 per day, and it uses these subcomponents at a rate of 12,500 per year (of 250 working days). Holding costs are $22 per item per year, and ordering costs are $32 per order.

a) What is the economic production quantity?

692.82 units (round your response to two decimal places).

b) How many production runs per year will be made?

production runs (round your response to two decimal places).
Answer #1

Solution:

Annual demand (D) = 12,500

Ordering cost (Co) = $32

Holding cost (H) = $22

Daily production (p) = 300 per day

Daily Demand (d) = Annual demand / Number of working days = 12,500 / 250 = 50 per day

(a) Economic production quantity (Q):

Q = SQRT [(2 x D x Co) / H x (1 - d/p)]

Q = SQRT [(2 x 12,500 x $32) / $22 x (1 - 50/300)]

Q = 208.89

Economic production quantity (Q) = 208.89 units

(b) Number of production runs:

Production runs = Annual demand / Economic production quantity

Production runs = 12,500 / 208.89

Production runs = 59.84